# spacing=nonfrench

### v1

@@ Abstract
Most adversarial threats in artificial intelligence (AI) target the computational behavior of models rather than the humans who rely on them. Yet modern AI systems increasingly operate within human decision loops, where users interpret and act on model recommendations. Large Language Models (LLMs) generate fluent natural-language explanations that shape how users perceive and trust AI outputs, revealing a new attack surface at the cognitive layer: the communication channel between AI and its users. We introduce adversarial explanation attacks (AEAs), where an attacker manipulates the framing of LLM-generated explanations to modulate human trust in incorrect outputs. We formalize this behavioral threat through the trust miscalibration gap, a metric that captures the difference in human trust between correct and incorrect outputs under adversarial explanations. By incorporating this gap, AEAs explore the daunting threats in which persuasive explanations reinforce users’ trust in incorrect predictions.
To characterize this threat, we conducted a controlled experiment (n = $205$), systematically varying four dimensions of explanation framing: reasoning mode, evidence type, communication style, and presentation format. Our findings show that users report nearly identical trust for adversarial and benign explanations, with adversarial explanations preserving the vast majority of benign trust despite being incorrect. The most vulnerable cases arise when AEAs closely resemble expert communication, combining authoritative evidence, neutral tone, and domain-appropriate reasoning. Vulnerability is highest on hard tasks, in fact-driven domains, and among participants who are less formally educated, younger, or highly trusting of AI. To our knowledge, this is the first systematic security study that treats explanations as an adversarial cognitive channel and quantifies their impact on human trust in AI-assisted decision making.
@@ 1 Introduction
Large Language Models (LLMs) epitomize this shift. Beyond producing predictions, they can generate additional fluent natural language explanations that accompany their output [breum2024persuasive, kang2024hallucinationtrust]. Explanations are designed to foster transparency and calibrate trust [miller2019explanation, ehsan2021human, liao2022questioning]. However, their persuasive tone and rhetorical coherence can powerfully shape human reasoning. Prior studies show that people often conflate confident articulation with correctness [jiang2023trust, xu2024trust]. This observation reframes the security boundary: an attacker who controls the presentation of reasoning may change user behavior. We argue that this form of behavioral manipulation represents a new class of adversarial threat that targets user cognition rather than model computation. Understanding such human-centric adversarial risks is essential to ensuring the robustness of AI-assisted decision-making systems.
In security terms, this behavioral threat extends beyond model manipulation into the human decision loop. Prior work [szegedy2014intriguing, goodfellow2015explaining, yuan2019adversarial] on adversarial machine learning largely targets the model itself, by altering inputs, parameters, or training data to change predictions. In contrast, adversarial explanation attacks (AEA) hold the model itself fixed and explore how the adversary manipulates the form and framing of LLM-generated explanations to bias users toward accepting incorrect outputs. This reveals a cognitive attack surface on the AI to human communication channel, which remains exposed even when the underlying model is robust. We quantify this vulnerability through a trust miscalibration gap: a behavioral analogue of adversarial success that captures how often users continue to trust incorrect model outputs under adversarial explanations. This metric provides a compact way to assess adversarial influence at the cognitive layer of AI systems. Although our empirical study centers on AI-assisted decision making, the same mechanism underlies a wide range of human-centric AI systems, from conversational copilots to autonomous decision interfaces, where explanation framing can shape user judgment [ehsan2021human, miller2019explanation, breum2024persuasive].
@@
RQ1: How to quantify the impact of adversarial explanations on human trust in AI-assisted decision making?
RQ2: Which explanation mechanisms and influencing factors moderate user trust in misleading explanations (e.g., task difficulty, domain, users’ education level)?
RQ3: How does user trust evolve over time under repeated exposure to adversarial explanations?
@@
Contributions. In this work, we treat explanations as a first-class security interface rather than a neutral transparency tool. We define adversarial explanation attacks that act entirely in the AI-to-human communication channel. Concretely, we (1) formalize the trust miscalibration gap as a behavioral analogue of adversarial success, (2) construct a four-dimensional space of LLM-generated explanation strategies expose a compositional attack surface, and (3) conduct a large-scale user study that quantifies which strategies, task contexts, and user groups produce the largest trust miscalibration gap with incorrect outputs. By shifting the adversarial lens from model robustness to cognitive robustness, our results argue that any security analysis of AI-assisted decision workflows must include the AI-to-human communication channel as an explicit attack surface.
@@ 3.2 Adversary Goals and Capabilities
We assume the adversary can influence the explanation-generation pipeline through practical attack vectors in modern LLM deployments, e.g., adversarial prompting that elicits persuasive yet misleading explanations [jones2023automated, zou2023universal], malicious fine-tuning or data poisoning alter reasoning structure or evidential framing [shen2023trigger, zou2023universal], and middleware hijacking that intercepts or rewrites explanation text at the API or plug-in layer [greshake2023more, liu2024jailbreak]. Representative scenarios include (1) a clinical decision support tool where prompt templates and explanation style are centrally configured and can be manipulated through prompt injection or malicious fine-tuning, and (2) an LLM-powered financial assistant where explanations are rendered in a web interface and middleware can intercept and rewrite explanation text but not model logits. In both settings, the adversary is constrained to generate explanations that remain syntactically valid and context-appropriate, and cannot alter the choice options shown to the user. Unlike classical attacks on model inputs or parameters, the behavioral adversary exploits cognitive vulnerabilities inside the human-AI communication loop.
@@ 4 Adversarial Explanation Framework
Building on the threat model in Section 3, we now operationalize how adversarial explanation attacks (AEAs) exploit the communication channel between an AI system and its human user. This section formalizes the behavioral attack surface of AI-generated explanations, framing it as a compositional space whose manipulation can modulate user trust and decision alignment.
4.1 Research Questions
Our goal is to empirically characterize how explanation framing shapes human trust in AI-assisted decision making, under both benign and adversarial conditions. To this end, we define three research questions that translate the abstract notion of communication integrity into measurable behavioral outcomes relevant to system security.
RQ1 (Adversarial Impact). Can adversarial explanations measurably distort user trust and decision alignment, even when the underlying AI outputs are incorrect?
RQ2 (Mechanisms and Moderators). Which explanation strategies, contextual conditions, and individual traits most strongly influence user susceptibility to persuasive framing?
RQ3 (Dynamic Trust Evolution). How does user trust recalibrate or erode with repeated exposure to misleading explanations over time?
Together, these questions extend adversarial learning beyond model-centric robustness to the human-AI decision loop. Rather than measuring classification error, we evaluate trust and trust miscalibration gap: how much users trust incorrect model outputs under adversarial explanations. The answers to these questions, derived from large-scale user studies described in Section 5, establish a behavioral foundation for reasoning about communication-space attacks on AI systems.
@@ 4.4 Operational Role in Adversarial Analysis
We now formalize how this compositional explanation space can be exploited by an adversary. For a given task $q$, an explanation framing strategy $s=(r^{\prime},v^{\prime},c^{\prime},p^{\prime})$ induces an explanation $e_{\text{A}}(q,s)$ generated by the LLM, while $e_{\text{B}}(q)$ denotes the benign explanation without framing strategy. Let $T(u,q,e)\in\mathbb{R}$ denote the user’s trial-level self-reported trust score for explanation $e$ on task $q$, measured on a 1177 Likert scale, which quantifies how much the user is willing to rely on the AI-provided answer given explanation $e$. We define the trust miscalibration gap as the change in user trust induced by adversarial explanation relative to the benign condition:
@@ 5.1 Unified Design and Experimental Setup
We test two conditions: (1) Benign, where the AI provides the correct answer with a valid explanation; and (2) Adversarial, where the AI justifies an incorrect answer using the same template. All other factors (task content, prompt format, explanation structure) remain identical, isolating the cognitive effect of adversarial framing from correctness or fluency.
@@ 5.2 Adversarial Explanation Generation and Control
Our adversarial explanation generation and control pipeline in Figure 1 ensures reproducibility, scalability, and adversarial realism through four sequential stages. Appendix A documents the strategy-guided explanation generation prompt (A.1), dataset reconstruction (A.4), and validation prompt (A.3).
@@
5.3 Human Study Procedure
@@ 6.1 Quantifying Adversarial Impact (RQ1)
RQ1: Whether adversarially framed explanations can measurably distort user trust and decision alignment, even when the underlying AI outputs are incorrect? We assess it by analyzing both (i) the cognitive sources that drive user trust and (ii) how trust scores shift under adversarial versus benign explanations.
@@ 6.1.1 Cognitive Sources of Human Trust
Figure 2: Proportion of trust cognitive sources under attacks and non-attacks.
@@ 6.2 Explanation Mechanisms and Influencing Factors (RQ2)
RQ2: Which explanation strategies, task circumstances, and user characteristics influence user vulnerability to persuasive explanation framing? The findings from RQ1 show that adversarial explanations sustain user trust even when outputs are incorrect. This motivates our next step: an analysis of why such misleading explanations remain persuasive and under what conditions they succeed. We answer RQ2 by analyzing user vulnerability at three levels: (1) explanation level mechanisms, (2) task level factors, and (3) user level traits.
@@ 6.3 Dynamic Trust Evolution (RQ3)
RQ3: How user trust recalibrates or erodes with repeated exposure to adversarial explanations over time? While prior analyses treat tasks as static, trust in AI is inherently dynamic: users may adapt, habituate, or desensitize to persuasive framing as they accumulate exposure to adversarial explanations.
@@ 7 Toward Defending Against Adversarial Explanations
Our empirical analysis across RQ1–RQ3 reveals a consistent vulnerability in AI-assisted decision making: persuasive explanations can modulate user trust. This shifts the adversarial threat surface from the model’s computational pipeline to the communication channel between AI and its users. When attackers shape explanation framing, they can influence human decisions even when the underlying model is correct. Securing AI systems, therefore, requires extending defenses beyond model robustness extending to explanation design, interfaces, and interaction protocols. Below, we outline a defense agenda informed by the vulnerabilities surfaced in our study.
Limit manipulative freedom degrees in explanations
@@
Adapt explanation style to contextual risk
@@ 9 Conclusion
As AI systems increasingly serve as both decision-makers and explainers, a new attack surface emerges: the communication channel through which AI explains its outputs to users. Our work introduces the adversarial explanation attacks (AEAs), where plausible yet misleading LLM-generated explanations distort user judgment and foster misplaced trust in incorrect outputs. We formalize a four-dimensional explanation strategy, including reasoning mode, evidence type, communication style, and presentation format, to systematically manipulate explanation framing and assess its impact on user trust. In a large-scale human study (n > 200), we find that users place nearly as much trust in adversarial explanations as in benign ones, with trust primarily attributed to the explanation itself. Moreover, user susceptibility varies by task context (e.g., difficulty, domain) and individual traits (e.g., education, AI familiarity), and repeated exposure to misleading explanations can gradually erode or recalibrate trust. These findings reveal critical cognitive vulnerabilities in human–AI interaction and highlight the need for future defenses that address not only model robustness but also users’ cognitive resilience.
@@
This work studies how AI-generated explanations influence human trust and decision-making under both benign and adversarial conditions. Since adversarial explanations directly shape human judgment, this research raises ethical concerns related to manipulation, trust miscalibration in human-AI interaction.
@@ A.2 Experiment Settings
A.3 LLM Quality Control and Configuration

### v2

@@ Abstract
Most adversarial threats in artificial intelligence (AI) target the computational behavior of models rather than the humans who rely on them. Yet modern AI systems increasingly operate within human decision loops, where users interpret and act on model recommendations. Large Language Models (LLMs) generate fluent natural-language explanations that shape how users perceive and trust AI outputs, revealing a new attack surface at the cognitive layer: the communication channel between AI and its users. We introduce adversarial explanation attacks (AEAs), where an attacker manipulates the framing of LLM-generated explanations to modulate human trust in incorrect outputs. We formalize this behavioral threat through the trust miscalibration gap, a metric that captures the difference in human trust between benign and adversarial explanations. Using this metric as a lens, we highlight a behavioral risk where persuasive explanation framing can preserve user trust even when the underlying AI prediction is wrong.
To characterize this threat, we conducted a human study with over $200$ participants, systematically varying four dimensions of explanation framing: reasoning mode, evidence type, communication style, and presentation format. Our findings show that users report nearly identical trust for adversarial and benign explanations, with adversarial explanations preserving the vast majority of benign trust despite being incorrect. The most vulnerable cases arise when AEAs closely resemble expert communication, combining authoritative evidence, neutral tone, and domain-appropriate reasoning. Vulnerability is highest on hard tasks, in fact-driven domains, and among participants who are less formally educated, younger, or highly trusting of AI. To the best of our knowledge, this is the first systematic security study that treats explanations as an adversarial cognitive channel and quantifies their impact on human trust in AI-assisted decision making.
@@ 1 Introduction
Large Language Models (LLMs) epitomize this shift. Beyond producing predictions, they can generate additional fluent natural language explanations that accompany their output [breum2024persuasive]. Explanations are designed to foster transparency and calibrate trust [miller2019explanation, liao2022questioning]. However, their persuasive tone and rhetorical coherence can powerfully shape human reasoning. Prior studies show that people often conflate confident articulation with correctness [moore2008trouble]. This observation reframes the security boundary: An attacker who controls the presentation of reasoning may change user behavior. We argue that this form of behavioral manipulation represents a new class of adversarial threat that targets user cognition rather than model computation. Understanding such human-centric adversarial risks is essential to ensuring the robustness of AI-assisted decision-making systems.
In security terms, this behavioral threat extends beyond model manipulation into the human decision loop. Prior work [szegedy2014intriguing, goodfellow2015explaining, yuan2019adversarial] on adversarial machine learning largely targets the model itself, by altering inputs, parameters, or training data to change predictions. In contrast, adversarial explanation attacks (AEAs) hold the model itself fixed and explore how the adversary manipulates the form and framing of LLM-generated explanations to bias users toward accepting incorrect outputs. This reveals a cognitive attack surface on the AI to human communication channel, which remains exposed even when the underlying model is robust. We quantify this vulnerability through a trust miscalibration gap: a behavioral analogue of adversarial success that captures how often users continue to trust incorrect model outputs under adversarial explanations. This metric provides a compact way to assess adversarial influence at the cognitive layer of AI systems. Although our empirical study centers on AI-assisted decision making, the same mechanism underlies a wide range of human-centric AI systems, from conversational copilots to autonomous decision interfaces, where explanation framing can shape user judgment [miller2019explanation, breum2024persuasive].
@@
RQ1 (Adversarial Impact): Can adversarial explanations measurably distort user trust and decision alignment, even when the underlying AI outputs are incorrect?
RQ2 (Mechanisms and Moderators): Which explanation strategies, contextual conditions, and individual traits most strongly influence user susceptibility to persuasive framing?
RQ3 (Dynamic Trust Evolution): How does user trust recalibrate or erode with repeated exposure to misleading explanations over time?
@@
Contributions. In this work, we treat explanations as a first-class security interface rather than a neutral transparency tool. We define adversarial explanation attacks that act entirely in the AI-to-human communication channel. Concretely, we (1) formalize the trust miscalibration gap as a behavioral analogue of adversarial success, (2) construct a four-dimensional space of LLM-generated explanation strategies that exposes a compositional attack surface, and (3) conduct a large-scale user study that quantifies which strategies, task contexts, and user groups produce the largest trust miscalibration gap with incorrect outputs. By shifting the adversarial lens from model robustness to cognitive robustness, our results argue that any security analysis of AI-assisted decision workflows must include the AI-to-human communication channel as an explicit attack surface.
@@ 3.2 Adversary Goals and Capabilities
We assume the adversary can influence the explanation-generation pipeline through practical attack vectors in modern LLM deployments, e.g., adversarial prompting that elicits persuasive yet misleading explanations [zou2023universal], malicious fine-tuning or data poisoning alter reasoning structure or evidential framing [zou2023universal, shu2023exploitability], and middleware hijacking that intercepts or rewrites explanation text at the API or plug-in layer [zhou2024trustworthiness]. Representative scenarios include (1) a clinical decision support tool where prompt templates and explanation style are centrally configured and can be manipulated through prompt injection or malicious fine-tuning, and (2) an LLM-powered financial assistant where explanations are rendered in a web interface and middleware can intercept and rewrite explanation text but not model logits. In both settings, the adversary is constrained to generate explanations that remain syntactically valid and context-appropriate, and cannot alter the choice options shown to the user. Unlike classical attacks on model inputs or parameters, the behavioral adversary exploits cognitive vulnerabilities inside the human-AI communication loop.
@@ 4 Adversarial Explanation Framework
Building on the threat model in Section 3, we now operationalize how adversarial explanation attacks (AEAs) exploit the communication channel between an AI system and its human user. This section formalizes the behavioral attack surface of AI-generated explanations, framing it as a compositional space whose manipulation can modulate user trust and decision alignment. The behavioral effects identified here are evaluated empirically through large-scale user studies described in Section 5.
@@ 4.3 Operational Role in Adversarial Analysis
We now formalize how this compositional explanation space can be exploited by an adversary. For a given task $q$, an explanation framing strategy $s=(r^{\prime},v^{\prime},c^{\prime},p^{\prime})$ induces an explanation $e_{\text{A}}(q,s)$ generated by the LLM, while $e_{\text{B}}(q)$ denotes the benign explanation without framing strategy. Rather than evaluating model-centric robustness or prediction accuracy, we focus on user trust and trust miscalibration gap: how much users trust incorrect model outputs under adversarial explanations. Let $T(u,q,e)\in\mathbb{R}$ denote the user’s trial-level self-reported trust score for explanation $e$ on task $q$, measured on a 1177 Likert scale, which quantifies how much the user is willing to rely on the AI-provided answer given explanation $e$. We define the trust miscalibration gap as the change in user trust induced by adversarial explanation relative to the benign condition:
@@ 5.1 Unified Design and Experimental Setup
We test two conditions: (1) Benign, where the AI provides the correct answer with a valid explanation; and (2) Adversarial, where the AI justifies an incorrect answer using the same template. All other factors (task content, prompt format, explanation structure) remain identical.
@@ 5.2 Adversarial Explanation Generation and Control
Our adversarial explanation generation and control pipeline in Figure 1 ensures reproducibility, scalability, and adversarial realism through four sequential stages. Appendix A documents the strategy-guided explanation generation prompt (A.1), dataset reconstruction (A.3), validation prompt (A.4).
@@
5.3 User Study Procedure
@@ 6.1 Quantifying Adversarial Impact (RQ1)
RQ1: Can adversarial explanations measurably distort user trust and decision alignment, even when the underlying AI outputs are incorrect? We assess it by analyzing (i) cognitive sources of trust and (ii) trust shifts under adversarial vs benign conditions.
@@ 6.1.1 Cognitive Sources of Human Trust
Figure 2: Proportion of trust cognitive sources under attacks and non-attacks. The majority of users attributed their trust to the explanation itself.
@@ 6.2 Explanation Mechanisms and Influencing Factors (RQ2)
RQ2: Which explanation strategies, contextual conditions, and individual traits most strongly influence user susceptibility to persuasive framing? The findings from RQ1 show that adversarial explanations sustain user trust even when outputs are incorrect. This motivates our next step: an analysis of why such misleading explanations remain persuasive and under what conditions they succeed. We answer RQ2 by analyzing user vulnerability at three levels: (1) explanation level mechanisms, (2) task level factors, and (3) user level traits.
@@ 6.3 Dynamic Trust Evolution (RQ3)
RQ3: How does user trust recalibrate or erode with repeated exposure to misleading explanations over time? While prior analyses treat tasks as static, trust in AI is inherently dynamic: users may adapt, habituate, or desensitize to persuasive framing as they accumulate exposure to adversarial explanations.
@@ 7 Toward Defending Against Adversarial Explanations
Our empirical analysis across RQ1–RQ3 reveals a consistent vulnerability in AI-assisted decision making: persuasive explanations can modulate user trust. This shifts the adversarial threat surface from the model’s computational pipeline to the communication channel between AI and its users. When attackers shape the framing of explanations, they can influence human decisions even when the underlying model is correct. Securing AI systems, therefore, requires extending defenses beyond model robustness to explanation design, interfaces, and interaction protocols. Below, we outline a defense agenda informed by the vulnerabilities surfaced in our study.
Limit the explanation framing strategies
@@
Risk-adaptive explanation policies
@@ 9 Conclusion
As AI systems increasingly serve as both decision-makers and explainers, a new attack surface emerges: the communication channel through which AI explains its outputs to users. Our work introduces the adversarial explanation attacks (AEAs), where plausible yet misleading LLM-generated explanations distort user judgment and foster misplaced trust in incorrect outputs. We formalize a four-dimensional explanation strategy, including reasoning mode, evidence type, communication style, and presentation format, to systematically manipulate explanation framing and assess its impact on user trust. By conducting a large-scale empirical user study ($n=205$), we find that users place nearly as much trust in adversarial explanations as in benign ones, with trust primarily attributed to the explanation itself. Moreover, user susceptibility varies by task context (e.g., difficulty, domain) and individual traits (e.g., education, age, initial trust in AI), and repeated exposure to misleading explanations can gradually erode or recalibrate trust. These findings reveal critical cognitive vulnerabilities in human–AI interaction and highlight the need for future defenses that address not only model robustness but also users’ cognitive resilience.
@@
This work studies how AI-generated explanations influence human trust and decision-making under both benign and adversarial conditions. Since adversarial explanations directly shape human judgment, this research raises ethical concerns related to manipulation and trust miscalibration in human-AI interaction.
@@ A.2 Experiment Settings
A.3 Dataset and Difficulty Calibration
A.4 LLM Quality Control and Configuration